Year Spot Rate 7% 8% 3 9% a) Given 84 = .04, S6 = .054, find i4,6 b) Given i1,3 = 5.6 and s3 = 6%, find sı
Answer #1

Ques 8). a). given 4 year spot rate s4 = 0.04

and 6 year spot rate s6 = 0.054

So, 2 year forward rate 4 years form now, i4,6 = (((1+s6)^6)/((1+s4)^4)^(1/2) - 1 = ((1.054^6)/(1.04^4)(1/2) - 1 = 8.26%

So, i4,6 = 8.26%

b). given 2 year forward rate 1 year from now i1,3 = 5.6%

3 year spot rate s3 = 6%

So, one year spot rate s1 = ((1+s3)^3)/((1+i1,3)^2) - 1 = (1.06^3)/(1.056^2) - 1 = 6.80%

So, s1 = 6.80%